OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 −x − y = −7 A. –y + 7 B. –y − 7 C. y + 7 D. y – 7 Is it D?

OpenStudy (anonymous):

Let's isolate x in the 2nd equation: -x-y=-7 -x=y-7 divide by -1: x=-y+7

OpenStudy (anonymous):

ohh, so its A. Okay, thanks for the help! :)